# KnowPhish: Large Language Models Meet Multimodal Knowledge Graphs for Enhancing Reference-Based Phishing Detection

### v1

@@ 3.1.3 Connection to Wikidata Knowledge Graph
To gain a deeper understanding on which Wikidata categories those phishing targets belong to, we perform searches within the $D_{2}$ dataset. Specifically, for each phishing target $b$, we search for the categories associated with it, denoted as $\mathcal{C}(b)=\{c|(b,\texttt{instance\_of},c)\in\mathcal{G}\}$. This process yields a collection of categories, $\mathcal{C}$, comprising the categories for all phishing targets in the APWG phishing dataset. In Figure 4, we present a visualization of the 30 most frequently occurring categories within $\mathcal{C}$. Our observation reveals that some categories represent specific industries, such as bank, postal service, and internet service provider, while others, like business, public company, and enterprise, convey more general semantics.
@@
Based on this observation, we curate two lists of categories: 1) Narrow Categories $\mathcal{C}_{n}$ such as ‘bank’ representing narrower industry segments, and 2) General Categories $\mathcal{C}_{g}$ such as ‘business’, which will be used in constructing our brand knowledge base KnowPhish. The detailed lists can be found in Table 9 and Table 10 in Appendix A.
@@ 3.2 Approach Overview
Figure 5: An overview of our automated pipeline for constructing our large scale multimodal BKB, KnowPhish. We first collect (a) all brands from certain high-value industries, and (b) only popular brands from general categories. Then, the knowledge acquisition and augmentation steps collect logos, domains and aliases for these brands.
Building upon the empirical insights in Section 3.1, we introduce KnowPhish, a large-scale multimodal BKB prioritizing both comprehensive coverage of potential phishing targets and detailed brand knowledge. As Figure 5 shows, KnowPhish is constructed using an automatic pipeline, which starts from 1) Brand Search based on industries via Wikidata knowledge graph $\mathcal{G}$ and proceeds with 2) Knowledge Acquisition and Augmentation to obtain relavant brand knowledge including logos, aliases, and domains. The complete construction algorithm is illustrated in Algorithm 2 in Appendix A.
@@ 3.3.1 Category-based Brand Search
$\mathcal{B}_{n}(c_{n})=\{b|(b,\texttt{instance\_of},c)\in\mathcal{G},c\in\{c_{n}\}\cup\mathcal{C}_{n}^{\prime}\}$ (1)
We introduce sub-categories here because the brands within a sub-category of a Narrow Category are also potential targets for phishing attackers. For example, National Bank of Costa Rica is categorized as state bank, a sub-category of bank, in Wikidata, although it unquestionably falls under the category of bank, and is a potential phishing target.
@@ 3.3.2 Popularity-based Brand Search
Although industry category is a key indicator of potential phishing targets, such information may not be accurately maintained in Wikidata for every brand. For example, Singtel, a Singaporean telco company, should logically fall into the telecommunication category. However, the only categories it has in Wikidata are ‘business’, ‘public company’, and ‘enterprise’. Hence, depending solely on Narrow Categories $\mathcal{C}_{n}$ is insufficient.
@@ 3.4.1 Knowledge Acquisition
where logo_image, official_website, and label222label is instantiated as ‘rdfs:label’ in SPARQL instead of a property relation. are the property relations in $\mathcal{G}$ that indicate the logos, URL of official website, and alternative names in different languages of an entity, respectively.
@@ 3.4.2 Knowledge Augmentation
Information maintained in Wikidata may be incomplete, particularly for the logos and the domains. Brands may employ multiple legitimate logo variants and domain variants in its online presence. When a phishing page contains a logo variant not present in the knowledge base, the phishing detector may fail to identify its brand, leading to false negatives. Similarly, if our detector examines a benign webpage with a legitimate domain that is not documented in Wikidata, a false positive alarm may be raised. Thus, further augmentation to $\mathcal{B}$ on the logos and domains is required to alleviate such false positives and false negatives.
Logo Variants.
To capture logo variants, we employ two methods. The first involves accessing the associated domain(s) of the brand and capturing the logo displayed on that webpage by a well-trained webpage layout detector[33], denoted $\textsf{DetectLogos}(b.domains)$. The second method utilizes Google Image Search[17]. We initiate a search query by combining the brand name with the term ‘logo’, then filter the results to include images with URLs matching the brand’s domain(s). In this way, we expand our logo collection beyond the Wikidata logo images:
$\begin{split}b.logos\leftarrow b.logos\ \cup\ \textsf{DetectLogos}(b.domains)\\ \ \cup\ \textsf{GoogleImageLogos}(b.name+\text{`logo'})\end{split}$ (2)
Domain Variants.
To acquire additional domain variants, we utilize the Tranco domain ranking list $\mathcal{D}$ and the Whois service[15]. Concretely, we run the Whois lookup on all the domains in KnowPhish and $\mathcal{D}$ to gather the Whois information for each of their domains. Then for each brand $b$, we expand the list of its legitimate domains by incorporating domains in $\mathcal{D}$ that share identical organization details with the original $b.domains$:
$\begin{split}b.domains\leftarrow b.domains\cup\{d|h_{whois}(d).org=\\ h_{whois}(b.domains).org,\ d\in\mathcal{D}\}\end{split}$ (3)
@@
Domain Propagation.
We further propose a method for propagating domain information among brand pairs that share subsidiary relationships, since the legitimate website of a brand may also display the logos of its subsidiary (or vice versa). For instance, ‘facebook.com’ can be seen as a domain variant for Meta. When visiting ‘facebook.com’, it is reasonable for a Meta logo to be present; but if we were not aware of this domain variant, we would classify it as having the brand intention of Meta, and thus a phishing attack, resulting in a false positive.
@@
Adapting to Evolving Phishing Targets.
New brands are continuously emerging as new potential phishing targets. Since KnowPhish is built in a fully automatic manner, one can simply handle such information obsolescence by regularly reconstructing KnowPhish to search for new potential phishing targets outside $\mathcal{B}$. Such regular updates allow the RBPD method to remain effective in countering attacks targeting these emerging brands.
@@ 4 KnowPhish Detector
Incorporating our multimodal BKB KnowPhish, we further propose KnowPhish Detector (KPD), a multimodal RBPD with multi-stage analysis. Figure 6 offers an overview of KPD, and Algorithm 1 further elaborates on its analysis steps. Specifically, for an input webpage $w$, KPD first leverages an LLM to generate a summary for $w$ using its HTML and URL. Then, the summary and the HTML are fed into a well-trained small language model to classify whether $w$ is a credential-requiring page (CRP). If $w$ is detected as a CRP, KPD will proceed to extract its brand intention from either the screenshot or the LLM summary. The extracted brand intention of $w$ are then used to retrieve a list of legitimate domains, which are compared with the domain of $w$ to decide whether $w$ is phishing or not. Notably, KPD features novel text-based modules to analyze implicit CRP and extract brand intention from logoless webpages, which will be discussed in Section 4.2 and Section 4.3, respectively.
@@ 4.1 LLM-based Webpage Summary
The LLM summary, as depicted in Figure 6, acts as crucial information for various subsequent tasks. To generate this summary, we first process the HTML by removing extraneous elements like JavaScript and CSS blocks. Subsequently, this processed HTML, along with the URL, will become the input of an LLM. We design a prompt template (Table 11 in Appendix B) with three in-context examples to understand the webpage from various aspects, such as what the brand intention of the webpage is, whether the webpage is a CRP, which elements in the HTML makes it a CRP, and the overall rationale for its CRP prediction. The generated summary will serve as an auxiliary text attribute of the webpage for the following CRP classification and text brand extraction tasks.
4.2 Text-based CRP Classification
As discussed in Section 2, phishing webpages always convey credential-requiring intentions. While a prior study has already developed an image-based CRP classifier[33], we empirically find that it leads to a significant number of false negatives during deployment by incorrectly classifying phishing webpages into non-CRP, in which most of these false negatives are implicit CRP. Figure 7 provides a comparison between explicit and implicit CRP, where explicit CRP directly shows credential-taking elements (e.g., the input fields of username and passwords), while implicit CRP only presents the elements (e.g., a login button) to redirect to a potential explicit CRP. While the credential-requiring intention is evident to page visitors, the CRP classifier from [33] cannot identify it because it exclusively focuses on identifying explicit CRPs.
(a) Explicit CRP
(b) Implicit CRP
Figure 7: Comparison between explicit and implicit CRP.
In response to this limitation, we propose using a small LM (XLM-RoBERTa[12]) to identify credential-requiring intention from text. Specifically, the small LM takes both the processed HTML and the LLM summary as input and outputs a binary prediction of whether the webpage is a CRP or not. This design of integrating the original text and its LLM summary as input to a smaller LM aims to benefit from the general-purpose reasoning and instruction following ability of an LLM, along with the trainability of a small LM [20]. Consequently, our text-based classifier can better comprehend the credential-requiring intention from webpage text, facilitating the detection of both explicit and implicit CRP. Webpages classified as non-CRP are regarded as benign; for those classified as CRP, we proceed to the brand extractor step.
4.3 Brand Extractor
Next, we aim to extract the brand intention of the webpage. Recall that our approach integrates with an existing RBPD which identifies brands visually through logos, which we call a logo brand extractor (LBE). Specifically, existing LBEs [32, 33] consists of a webpage layout detector to locate a logo from the screenshot, and a logo matcher to match that logo to a brand in the BKB. We find that existing RBPDs encounter limitations in identifying brand intention when 1) the logo displayed in the webpage differs from the logos stored in the BKB, or 2) the logo cannot even be detected (e.g., Figure 1).
In situations where the LBE fails to detect a logo on the webpage or cannot find a matching logo, the TBE is activated to extract brand intention, leading to higher recall.
4.4 Domain Check
@@
[…]
@@ 5.1 Datasets
We utilize two datasets for our main phishing detection experiments. 1) TR-OP: A manually labeled and balanced dataset where the benign samples are collected from Tranco[44] and the phishing samples are obtained from OpenPhish[40]. The phishing samples were crawled and validated within 6 months from July to December 2023 and covered 440 unique phishing targets. Note that the phishing samples here are different from the APWG phishing dataset $D_{2}$ from Section 3.1. 2) SG-SCAN: An unlabelled dataset where the samples come from Singapore’s local webpage traffic. We randomly sample 10k webpages dating from mid-August 2023 to mid-January 2024. It is used to evaluate the phishing detection approaches in the local setting. Table 1 offers an overview of both datasets.
@@
In addition, we also manually extracted and labelled 2555 samples to train a XLM-RoBERTa [12], our text-based CRP classifier. This dataset contains 1094 phishing samples from $D_{2}$ and 1461 benign samples from Alexa Ranking [4]. The 1094 phishing samples are all CRP. Among the 1461 benign samples, 1297 are non-CRP, while the remaining 164 are CRP. After combining these samples, they are divided into 0.8/0.1/0.1 train/valid/test splits.
@@ 5.2 Baselines
We select two state-of-the-art approaches, Phishpedia[32] and PhishIntention[33], together with our proposed KPD as the RBPD backbones. Both Phishpedia and PhishIntention can be either equipped with their original reference list (containing 277 brands), DynaPhish[34], or our proposed KnowPhish, as the BKB used for phishing detection. KPD will only be equipped with KnowPhish due to the requirement of alias information. For fair comparison, both KnowPhish and DynaPhish will construct their knowledge from an empty BKB, since one can always improve the performance of both knowledge expansion approaches by manually adding well-inspected brand knowledge.
We assume a static environment in all our experiments (i.e., the only data available on a webpage is its URL, screenshot, and HTML). In this case, the dynamic analysis module in PhishIntention and the webpage interaction module in DynaPhish will be disabled. Further details on the implementation can be found in Appendix C.
@@ 5.3 RQ1: Effectiveness and Efficiency
Detector BKB ACCnormal-$\uparrow$ F1normal-$\uparrow$ Precisionnormal-$\uparrow$ Recallnormal-$\uparrow$ Timenormal-$\downarrow$
Original
68.17
53.54
99.08
36.68
0.22s
Phishpedia
DynaPhish
66.40
52.52
89.50
37.16
10.92s
KnowPhish
85.79
83.67
98.27
72.80
0.22s
Original
64.97
46.15
99.73
30.02
0.24s
PhishIntention
DynaPhish
62.51
41.16
95.62
26.22
10.67s
KnowPhish
77.84
71.60
99.67
55.84
0.26s
KPD
KnowPhish
92.49 92.05
97.84
86.90
2.02s
Table 2: Phishing detection performance of different RBPDs on TR-OP dataset. The metric ‘Time’ indicates the average processing time per sample, while the remaining metrics are presented in percentages. $\uparrow$ means higher is better while $\downarrow$ refers to the opposite.
@@
(a) Logos in phishing webpages
(b) Logos in legitimate webpages
Figure 8: Examples of logo variants on phishing samples and legitimate homepages where DynaPhish fails to identify.
@@
  • KnowPhish substantially boosts the F1 score of Phishpedia by 30% and PhishIntention by 25%, and also increases their recall by 25%, with only marginal impacts on precision, compared to other BKB baselines. The primary factor contributing to the superior performance of KnowPhish over DynaPhish is that KnowPhish contains more logo variants. In contrast, DynaPhish can only identify the logos identical to those displayed on official webpages (a few examples are provided in Figure 8).
  • KPD provides the highest F1 score of 92.02%, and recall of 86.90%, substantially outperforming other approaches. KPD benefits from the alias information from KnowPhish, allowing it to detect logo-less phishing pages. Consequently, KPD identifies more phishing targets than DynaPhish, as shown in Figure 9.
  • KnowPhish achieves runtime efficiency by decoupling the BKB construction from phishing detectors. Unlike DynaPhish which requires crawling additional webpages during deployment, KnowPhish identifies potential phishing targets locally, leading to about 50 times lower running time when integrated with Phishpedia and PhishIntention. Even when equipped with KPD with additional LLM query overhead, KnowPhish remains 5 times faster than DynaPhish with image-based RBPDs.
Overall, KnowPhish and KPD increase the coverage of brands and their logos, and allow us to detect logo-less pages, substantially improving detection performance, while maintaining acceptable runtime overhead.

### v2

@@ 4.1.3 Connection to Wikidata Knowledge Graph
To gain a deeper understanding of which Wikidata categories those phishing targets belong to, we perform searches within the $D_{2}$ dataset. Specifically, for each phishing target $b$, we search for the categories associated with it, denoted as $\mathcal{C}(b)=\{c|(b,\texttt{instance\_of},c)\in\mathcal{G}\}$. This process yields a collection of categories, $\mathcal{C}$, comprising the categories for all phishing targets in the APWG phishing dataset. In Figure 4, we present a visualization of the 30 most frequently occurring categories within $\mathcal{C}$. Our observation reveals that some categories represent specific industries, such as bank, postal service, and internet service provider, while others, like business, public company, and enterprise, convey more general semantics.
Based on this observation, we curate two lists of categories: 1) Narrow Categories $\mathcal{C}_{n}$ such as ‘bank’ representing narrower industry segments, and 2) General Categories $\mathcal{C}_{g}$ such as ‘business’, which will be used in constructing our brand knowledge base KnowPhish. The detailed lists can be found in our supplementary materials[26].
@@ 4.2 Approach Overview
Figure 5: An overview of our automated pipeline for constructing our large-scale multimodal BKB, KnowPhish. We first collect (a) all brands from certain high-value industries, and (b) only popular brands from general categories. Then, the knowledge acquisition and augmentation steps collect logos, domains, and aliases for these brands.
Building upon the empirical insights in Section 4.1, we introduce KnowPhish, a large-scale multimodal BKB prioritizing both comprehensive coverage of potential phishing targets and detailed brand knowledge. As Figure 5 shows, KnowPhish is constructed using an automatic pipeline, which starts from 1) Brand Search based on industries via Wikidata knowledge graph $\mathcal{G}$ and proceeds with 2) Knowledge Acquisition and Augmentation to obtain relevant brand knowledge including logos, aliases, and domains. The complete construction algorithm is illustrated in our supplementary materials[26].
@@ 4.3.1 Category-based Brand Search
$\mathcal{B}_{n}(c_{n})=\{b|(b,\texttt{instance\_of},c)\in\mathcal{G},c\in\{c_{n}\}\cup\mathcal{C}_{n}^{\prime}\}$
We introduce sub-categories here because the brands within a sub-category of a Narrow Category are also potential targets for phishing attackers. For example, National Bank of Costa Rica is only categorized as ‘state bank’, a sub-category of ‘bank’, in Wikidata. Despite this, it unquestionably falls under the category of ‘bank’, and is a potential phishing target.
@@ 4.3.2 Popularity-based Brand Search
Although the industry category is a key indicator of potential phishing targets, such information may not be accurately maintained in Wikidata for every brand. For example, Singtel, a Singaporean telco company, should logically fall into the telecommunication category. However, the only categories it has in Wikidata are ‘business’, ‘public company’, and ‘enterprise’. Hence, depending solely on Narrow Categories $\mathcal{C}_{n}$ is insufficient.
@@ 4.4.1 Knowledge Acquisition
where logo_image, official_website, and label are the property relations in $\mathcal{G}$ that indicate the logos, URL of the official website, and alternative names in different languages of an entity, respectively.
@@ 4.4.2 Knowledge Augmentation
Information maintained in Wikidata may be incomplete, particularly for the logos and the domains. Brands may employ multiple legitimate logo variants and domain variants in their online presence. When a phishing page contains a logo variant not present in the knowledge base, the phishing detector may fail to identify its brand, leading to false negatives. Similarly, if our detector examines a benign webpage with a legitimate domain that is not documented in Wikidata, a false positive alarm may be raised. Thus, further augmentation to $\mathcal{B}$ on the logos and domains is required to alleviate such false positives and false negatives.
Logo Variants.  To capture logo variants, we employ two methods. The first involves accessing the associated domain(s) of the brand and capturing the logo displayed on that webpage by a well-trained webpage layout detector[34], denoted $\textsf{DetectLogos}(b.domains)$. The second method utilizes Google Image Search[18]. We initiate a search query by combining the brand name with the term ‘logo’, then filter the results to include images with URLs matching the brand’s domain(s). In this way, we expand our logo collection beyond the Wikidata logo images:
$\begin{split}b.logos\leftarrow b.logos\ \cup\ \textsf{DetectLogos}(b.domains)\\ \ \cup\ \textsf{GoogleImageLogos}(b.name+\text{`logo'})\end{split}$
Domain Variants.  To acquire additional domain variants, we utilize the Tranco domain ranking list $\mathcal{D}$ and the Whois service[16]. Concretely, we run the Whois lookup on all the domains in KnowPhish and $\mathcal{D}$ to gather the Whois information for each of their domains. Then for each brand $b$, we expand the list of its legitimate domains by incorporating domains in $\mathcal{D}$ that share identical organization details with the original $b.domains$:
$\begin{split}b.domains\leftarrow b.domains\cup\{d|h_{whois}(d).org=\\ h_{whois}(b.domains).org,\ d\in\mathcal{D}\}\end{split}$
@@
Domain Propagation.  We further propose a method for propagating domain information among brand pairs that share subsidiary relationships, since the legitimate website of a brand may also display the logos of its subsidiary (or vice versa). For instance, ‘facebook.com’ can be seen as a domain variant for Meta. When visiting ‘facebook.com’, it is reasonable for a Meta logo to be present; but if we were not aware of this domain variant, we would classify it as having the brand intention of Meta, and thus a phishing attack, resulting in a false positive.
@@
Adapting to Evolving Phishing Targets. New brands are continuously emerging as potential phishing targets. Since KnowPhish is built in a fully automatic manner, one can simply handle such information obsolescence by regularly reconstructing KnowPhish to search for new potential phishing targets outside $\mathcal{B}$. Such regular updates allow the RBPD method to remain effective in countering attacks targeting these emerging brands.
Adversarial Injection  Despite Wikidata being well-maintained, there remains a risk of attackers inserting phishing URLs into its database. To mitigate this, we verify URLs against existing phishing blacklists (e.g., Google Safe Browsing[47]) before adding them to our BKB.
@@ 5 KnowPhish Detector
Incorporating our multimodal BKB KnowPhish, we further propose KnowPhish Detector (KPD), a multimodal RBPD with multi-stage analysis. Figure 6 offers an overview of KPD, and Algorithm 1 further elaborates on its analysis steps. Specifically, for an input webpage $w$, KPD first leverages an LLM to generate a summary for $w$ using its HTML and URL. Then, the summary and the HTML are fed into a well-trained small language model to classify whether $w$ is a credential-requiring page (CRP). If $w$ is detected as a CRP, KPD will proceed to extract its brand intention from either the screenshot or the LLM summary. The extracted brand intention of $w$ is then used to retrieve a list of legitimate domains, which are compared with the domain of $w$ to decide whether $w$ is phishing or not. Notably, KPD features novel text-based modules to analyze implicit CRP and extract brand intention from logo-less webpages, which will be discussed in Section 5.2 and Section 5.3, respectively.
@@ 5.1 LLM-based Webpage Summary
The LLM summary, as depicted in Figure 6, acts as crucial information for various subsequent tasks. To generate this summary, we first process the HTML by removing extraneous elements like JavaScript and CSS blocks. Subsequently, this processed HTML, along with the URL, will become the input of an LLM. We design a prompt template (details provided in [26]) with three in-context examples to understand the webpage from various aspects, such as what the brand intention of the webpage is, whether the webpage is a CRP, which elements in the HTML makes it a CRP, and the overall rationale for its CRP prediction. The generated summary will serve as an auxiliary text attribute of the webpage for the following CRP classification and text brand extraction tasks.
1
2
@@
10       $b_{t}\leftarrow\textsf{AliasMap}(\textsf{TextBrand}(\textit{s}))$
@@
5.2 Text-based CRP Classification
As discussed in Section 3, phishing webpages always convey credential-requiring intentions. While a prior study has already developed an image-based CRP classifier[34], we empirically find that it leads to a significant number of false negatives during deployment by incorrectly classifying phishing webpages into non-CRP, in which most of these false negatives are implicit CRP. Figure 7 provides a comparison between explicit and implicit CRP, where explicit CRP directly shows credential-taking elements (e.g., the input fields of username and passwords), whereas implicit CRP only presents the elements (e.g., a login button) to redirect to a potential explicit CRP. While the credential-requiring intention is evident to page visitors, the CRP classifier from [34] cannot identify it because it exclusively focuses on identifying explicit CRPs.
(a) Explicit CRP
(b) Implicit CRP
Figure 7: Comparison between explicit and implicit CRP.
In response to this limitation, we propose using a small multilingual LM (XLM-RoBERTa[13]) to identify credential-requiring intention from text. Specifically, the small LM takes both the processed HTML and the LLM summary as input and outputs a binary prediction of whether the webpage is a CRP or not. This design of integrating the original text and its LLM summary as input to a smaller LM aims to benefit from the general-purpose reasoning and instruction following ability of an LLM, along with the trainability of a small LM [21]. Consequently, our text-based classifier can better comprehend the credential-requiring intention from webpage text, facilitating the detection of both explicit and implicit CRP. Webpages classified as non-CRP are regarded as benign; for those classified as CRP, we proceed to the brand extractor step.
5.3 Brand Extractor
Next, we aim to extract the brand intention of the webpage. Recall that our approach integrates with an existing RBPD which identifies brands visually through logos, which we call a logo brand extractor (LBE). Specifically, existing LBEs [33, 34] consists of a webpage layout detector to locate a logo from the screenshot, and a logo matcher to match that logo to a brand in the BKB. We find that existing RBPDs encounter limitations in identifying brand intention when 1) the logo displayed on the webpage differs from the logos stored in the BKB, or 2) the logo cannot even be detected.
In situations where the LBE fails to detect a logo on the webpage or cannot find a matching logo, the TBE is activated to extract brand intention, leading to higher recall. TBE is similar to the counterfactual analysis module [35] in its ability to detect logo-less pages. However, the counterfactual analysis focuses on interacting with the webpage, e.g. by verifying credentials or redirection, while TBE focuses on detecting brand intention through the textual information on the webpage.
5.4 Domain Check
@@ 6.1 Datasets
We utilize two datasets for our main phishing detection experiments. 1) TR-OP: A manually labeled and balanced dataset where the benign samples are randomly collected from Tranco[46] top 50k domains and the phishing samples are obtained from OpenPhish[42]. The phishing samples were crawled and validated within 6 months from July to December 2023, covering 440 unique phishing targets. Note that the phishing samples here are different from $D_{1}$ and $D_{2}$ discussed in Section 4.1. 2) SG-SCAN: An unlabelled dataset with samples from Singapore’s local webpage traffic. We randomly sample 10k webpages dating from mid-August 2023 to mid-January 2024. It is used to evaluate the phishing detection approaches in the local context. Table 1 offers an overview of both datasets.
@@
In addition, we also manually extracted and labeled 2555 samples to train an XLM-RoBERTa [13], our text-based CRP classifier. This dataset contains 1094 phishing samples from $D_{2}$ and 1461 benign samples from Alexa Ranking [4]. The 1094 phishing samples are all CRP. Among the 1461 benign samples, 1297 are non-CRP, while the remaining 164 are CRP. After combining these samples, they are divided into 0.8/0.1/0.1 train/valid/test splits.
@@ 6.2 Baselines
We select two state-of-the-art approaches, Phishpedia[33] and PhishIntention[34], together with our proposed KPD as the RBPD backbones. Both Phishpedia and PhishIntention can be either equipped with their original reference list (containing 277 brands), DynaPhish[35], or our proposed KnowPhish, as the BKB used for phishing detection. Due to the requirement of alias information, KPD will be only equipped with KnowPhish or an extended version of DynaPhish. In this extended DynaPhish, the extracted brand name will be used as the only alias of each new brand. For fair comparison, both KnowPhish and DynaPhish will construct their knowledge from an empty BKB, since one can always improve the performance of both knowledge expansion approaches by manually adding well-inspected brand knowledge.
We assume a static environment in all our experiments (i.e., the only data available on a webpage is its URL, screenshot, and HTML). In this case, the dynamic analysis module in PhishIntention and the webpage interaction module in DynaPhish will be disabled. Further details on the implementation can be found in Appendix A.1.
@@ 6.3 RQ1: Effectiveness and Efficiency
  • KnowPhish substantially boosts the F1 score of Phishpedia by 25% and PhishIntention by 20%, and also increases their recall by 32% and 22%, with only marginal impacts on precision, compared to other BKB baselines. The primary factor contributing to the superior performance of KnowPhish over DynaPhish is that it already encompasses most phishing targets and their logo variants. In contrast, DynaPhish suffers from the logo-matching constraint required to build brand knowledge, causing many false negatives as illustrated in Figure 1.
  • KPD+KnowPhish provides the highest F1 score of 92.05%, and recall of 86.90%, substantially outperforming other approaches. KPD benefits from the rich alias information from KnowPhish, allowing it to detect logo-less phishing pages through analysis of HTML and URL. Consequently, KPD identifies more phishing targets than DynaPhish, as shown in Figure 8.
  • KnowPhish achieves better runtime efficiency than DynaPhish by decoupling the BKB construction from phishing detectors. Unlike DynaPhish which requires crawling additional webpages during deployment, KnowPhish identifies potential phishing targets locally, leading to about 50 times lower running time when integrated with Phishpedia and PhishIntention. Even when equipped with KPD with additional LLM query overhead, KnowPhish remains 6 times faster than DynaPhish.
[…]
Original 69.91 57.17 99.16 40.16 0.25s
Phishpedia DynaPhish 66.40 52.52 89.50 37.16 10.92s
KnowPhish 85.79 83.67 98.27 72.80 0.22s
Original 66.62 49.96 99.76 33.32 0.28s
PhishIntention DynaPhish 62.51 41.16 95.62 26.22 10.67s
KnowPhish 77.84 71.60 99.67 55.84 0.26s
KPD DynaPhish 76.10 69.71 95.16 55.00 12.18s
KnowPhish 92.49 92.05 97.84 86.90 2.02s
Table 2: Phishing detection performance of different RBPDs on TR-OP dataset. The metric ‘Time’ indicates the average inference time per sample, while the remaining metrics are presented in percentages. $\uparrow$ means higher is better while $\downarrow$ refers to the opposite.